# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 46                   Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
  • When $\mu=3$, let us consider two subcases. If $c_{i}(e_{2})=b_{i}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}=c_{i}(e_{1})$ and $c_{i}(f_{2})=b_{i}=c_{i}(e_{2})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=a_{j}=c_{j}(f_{1})$ and $c_{j}(e_{2})\leq c_{j}(f_{2})=b_{i}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $c_{i}(e_{2})=a_{i}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{2})=a_{i}=c_{i}(e_{1})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda=1$ and continue executing this algorithm.
@@
  • When $\mu=4$, agent $j$ picks it, For agent $i$, her bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{2})\leq c_{j}(e_{1})=b_{j}$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda=1$ and continue executing this algorithm.
@@
In Algorithm 8, we follow a similar way of offline envy graph procedures to allocate items. The precondition of Algorithm 8 is that there is no envy cycle between two agents. Initially, the empty allocation is EF1 without an envy cycle. When there is no envy between two agents, we allocate the arriving item to one agent according to different conditions. When there exists an envy between two agents, we have two cases. One is that we allocate it to the unenvied agent when the envy cycle does not appear. The other is that if we allocate it to the unenvied agent, the envy cycle will appear, and we terminate Algorithm 8. Similar to the allocation of goods, we need to ensure that there is no envy cycle between two agents. Because we can not revoke the previous allocation in the online setting, Algorithms 9 and 10 help us eliminate the possible envy cycle. When the allocation $A$ satisfies the precondition and is EF1, we will continue to execute Algorithm 8. By the induction, it is not hard to see that the output allocation of Algorithm 8 is always EF1 without an envy cycle. Combining Lemmas 4, 5, and the description of Algorithm 8, we can conclude that the output complete allocation is $2$-EF1. For MMS, by the implication between EF1 and MMS [22], i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{n\alpha+n-1}{n-1+\alpha}$-MMS, where $n$ is the number of agents, it is easy to see that the output allocation is $\frac{5}{3}$-MMS. ∎
@@
For two agents with additive bi-valued cost functions, no deterministic online algorithm can compute a complete allocation that is $(2-\epsilon)$-EF1 or $(\frac{3}{2}-\epsilon)$-MMS for any $\epsilon>0$.
@@
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
@@
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ $1$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
@@
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
@@
Table 7: The impossibility results of EF1 for bi-valued cost functions for two agents.
$t=1$ $t=2$
agent $1$ $1$ $1$
agent 2 $1$ $1$
@@
Table 8: The impossibility results of MMS for additive bi-valued cost functions for two agents.
Consider an online allocation instance with two agents $\{1,2\}$. For any agent $i\in\{1,2\}$ and any item $e$, $c_{i}(e)\in\{1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1. The cost of each arrived item is shown in Table 9. To guarantee the completeness of chores allocation, we should guarantee that each arriving item is allocated to one agent.
The impossibility of EF1.
When $t=1$, the arriving item has a cost of 1 for both agents. Without loss of generality, assume that agent 1 picks it. When $t=2$, the item that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. To avoid the approximation ratio of EF1 being infinity, it is allocated to agent 2. Then, it is easy to check that the allocation is EF1 and MMS. In round $t=3$, the item that has the cost of $\frac{1}{\epsilon}$ for both agents arrives. Then, let us consider two cases.
@@
For the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation that is $\alpha$-EF1 or $\beta$-MMS, and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$, even for two agents with additive bi-valued cost functions.
@@
Consider an online allocation instance with two agents $\{1,2\}$, $c_{1}\in\{1,2\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+\epsilon}{\epsilon+1}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha\geq 1$ or $\beta$-MMS for any $\beta<2$, and $\gamma$-min-USC for any given $\gamma\geq 1$ when two agents have the bi-valued cost functions. ∎
@@ 4.3.2 The Combination with Additive Binary Costs
Next, we consider the setting where one agent (agent $n$) has the additive bi-valued cost function, and the remaining agents have additive binary cost functions and propose the Adapted-Chores-Picking Algorithm (Algorithm 11) to compute a complete allocation that satisfies EF1 and MMS. The intuition of this algorithm is similar to the goods setting. In Table 9, we illustrate how our algorithm works for three agents. Regarding efficiency, we show that EF1 or MMS is incompatible with USC in this setting.
-Input: An instance $(T,N,(c_{i})_{i\in N})$ with $(n-1)$ additive binary cost functions $(c_{1}.\ldots,c_{n-1})$ and one additive bi-valued cost function $(c_{n})$.
Output: A complete EF1 and MMS allocation $A$
@@
2Let $A=(\emptyset,\ldots,\emptyset)$;
@@ 25                         $A_{n}=A_{n}\cup\{e\}$;
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ldots
agent $1$ $1$ 00 $1$ 1 $1$ldots
agent $2$ $1$ 1 1 $1$ 1 $\ldots$
agent 3 $a_{3}$ $a_{3}$ $a_{3}$ $b_{3}$ $b_{3}$ldots
@@
For the deterministic allocation of indivisible chores for $(n-1)$ agents with additive binary cost functions and one agent with the additive bi-valued cost function, the Adapted-Chores-Picking Algorithm (Algorithm 11) computes a complete allocation that satisfies EF1 and MMS.
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 11.
Firstly, we show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after the round $k$, the allocation $A^{k}$ is EF1. We show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is also EF1.
@@
  • If $c_{n}(e_{k+1})=a_{n}$, then let us consider two subcases. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\geq|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is still EF1. For agent $n$, her bundle is EF1 since she has the priority of picking the item with the cost of $a_{n}$ and the number of items in her bundle is no more than that of items with non-zero costs in agent $i_{min}$’s bundle. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|<|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the binary cost valuation and the lowest number of items with non-zero costs.
  • If $v_{n}(e_{k+1})=b_{n}$, the let us consider two subcases. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\leq|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the binary cost valuation and the lowest number of items with non-zero costs. For any agent $i\in N\setminus\{i_{min}\}$, her bundle remains untouched and is still EF1. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|>|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is trivially EF1. For agent $n$, her bundle is EF since she has the priority of picking the item with the cost of $a_{n}$, and in this case, the extra one item in another agent’s bundle must have the cost of $b_{n}$ for agent $n$.
Combining the above cases, it can be seen that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is still EF1.
Next, we will show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is MMS. For arbitrary agent $i\in N$, if her bundle is EF, we are done. If not, then let us consider two cases. When $i\in[n-1]$, she has a binary valuation function, and by our algorithm description, the number of items with non-zero costs in her bundle is at most one more than that of the other agent’s bundle. Thus, it is not hard to see that her bundle is MMS. When $i=n$, currently, the number of items in her bundle is one more than that of the other agent’s bundle, and the extra item has the cost of $a_{n}$. After removing this item, the cost of her bundle is no more than that of the other agent’s bundle in her perspective, so we can conclude that her bundle is MMS. ∎
@@
For the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation that is $\alpha$-EF1 or MMS, and $\beta$-min-USC for any $\alpha,\beta\geq 1$, even for two agents where one agent has the additive binary cost function and the other agent has the additive bi-valued cost function.
@@
Without loss of generality, we assume that $c_{1}\in\{0,1\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+1}{2}=\frac{1}{2\epsilon^{2}}+\frac{1}{2}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm outputs a complete allocation that is $\alpha$-EF1 or MMS in the setting of binary and bi-valued cost functions, and guarantees $\beta$-min-USC for any given $\alpha,\beta\geq 1$. ∎

### v2

@@ 31    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it. Then, we set $\lambda_{2}=1$ and $\mu_{2}=1$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=b_{j}\leq c_{j}(f_{2})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=a_{j}$, agent $i$ picks it. Then, we set $\lambda_{2}=2$ and $\mu_{2}=2$. For agent $i$, her new bundle is still 2-EF1 since we have $\frac{c_{i}(A_{i}\setminus\{f\}\cup\{e_{1}\})}{c_{i}(A_{j})}\leq\frac{c_{i}(A_{j})+c_{i}(e_{1})}{c_{i}(A_{j})}\leq 2$ for some $f\in\{f_{1},f_{2}\}$ where the inequality follows from the fact that $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}$ and $|A_{j}|\geq 2$. Thus, this allocation is 2-EF1.
It can be seen that in the last two cases, the CDBC algorithm continues to be executed.
$\lambda_{2}=2$. When the item $e_{2}$ arrives at $\lambda_{2}=2$, let us consider two cases about $\mu_{2}$.
@@
  • When $\mu_{2}=1$, we need to consider two subcases. If $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$ and $c_{j}(e_{2})\leq c_{i}(e_{1})=b_{i}$. For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=b_{j}=c_{j}(e_{2})$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda_{2}=1$ and $\mu_{2}=0$, implying that the CDBC algorithm continues to be executed. If $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=b_{j}=c_{j}(f_{2})$ and $c_{j}(e_{2})=a_{j}=c_{j}(f_{1})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$, implying that the CEGP algorithm is ready to be executed.
  • When $\mu_{2}=1$, agent $j$ picks it, For agent $i$, her bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{2})\leq c_{j}(e_{1})=b_{j}$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda_{2}=1$ and $\mu_{2}=0$, implying that the CDBC algorithm continues to be executed.
@@
In the main body of the CAGEP algorithm 9, it can be seen that for different values of the algorithm indicator $\alpha$, we need to choose the corresponding algorithm to execute. By Lemmas 8, 9, and 10, it can be concluded that the output allocation is complete and $2$-EF1. Regarding MMS, by the implication between EF1 and MMS Sun et al. (2021), i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{n\alpha+n-1}{n-1+\alpha}$-MMS, where $n$ is the number of agents, it is easy to see that the output allocation is $\frac{5}{3}$-MMS. ∎
Next, we complement our approximation results by providing the lower bounds below.
@@
For two agents with additive personalized bi-valued cost functions, no deterministic online algorithm can compute a complete allocation that is $(2-\epsilon)$-EF1 or $(\frac{3}{2}-\epsilon)$-MMS for any $\epsilon>0$.
@@
$e_{1}$ $e_{2}$ $e_{3}$
agent 11 11 11 $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
@@
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 11 11 $\frac{1}{\epsilon}$ 11
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
@@
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
agent 11 11 11 $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
@@
Table 10: The impossibility results of EF1 for additive personalized bi-valued cost functions for two agents.
$e_{1}$ $e_{2}$
agent 11 11 11
agent 2 11 11
@@
Table 11: The impossibility results of MMS for additive personalized bi-valued cost functions for two agents.
Let $a_{1}=a_{2}=1$ and $b_{1}=b_{2}=\frac{1}{\epsilon}$ where $0<\epsilon<1$. Consider the following online instance for $n=2$, where the cost of each item is shown in Table 12. To guarantee the completeness of the chores allocation, we should guarantee that each arriving item is allocated to one agent.
EF1.
When $t=1$, the arriving item has the cost of 1 for both agents. Without loss of generality, assume that agent 1 picks it. When $t=2$, the item that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. To avoid the approximation ratio of EF1 being infinity, it is allocated to agent 2. Then, it is easy to check that the allocation is EF1 and MMS. In round $t=3$, the item that has the cost of $\frac{1}{\epsilon}$ for both agents arrives. Then, let us consider two cases.
@@
At last, we examine the compatibility between approximate EF1 or MMS and USC, and provide the following negative results.
@@
Given any $\alpha,\gamma\geq 1$ and $\beta<2$, for the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation satisfying
  • $\alpha$-EF1 and $\gamma$-min-USC,
  • or $\beta$-MMS and $\gamma$-min-USC,
even for two agents with additive personalized bi-valued cost functions.
@@
Without loss of generality, we assume that $c_{1}\in\{1,2\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+\epsilon}{\epsilon+1}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha\geq 1$ or $\beta$-MMS for any $\beta<2$, and $\gamma$-min-USC for any given $\gamma\geq 1$ when two agents have the additive personalized bi-valued cost functions. ∎
@@ 4.3.2 The Combination with Additive Binary Costs
Next, we consider the setting where one agent (agent $n$) has the additive personalized bi-valued cost function, and the remaining agents have additive binary cost functions and propose the Adapted-Chores-Picking Algorithm (Algorithm 13) to compute a complete allocation that satisfies EF1 and MMS. The intuition of this algorithm is similar to the goods setting. In Table 12, we illustrate how our algorithm works for three agents. Regarding efficiency, we show that EF1 or MMS is incompatible with USC in this setting.
Input: An instance $(T,N,(c_{i})_{i\in N})$ with (n1)(n-1) additive binary cost functions $(c_{1}.\ldots,c_{n-1})$ and one additive personalized bi-valued cost function $(c_{n})$.
Output: A complete EF1 and MMS allocation $\bf A$}
@@
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
@@ 25          $A_{n}=A_{n}\cup\{e\}$;
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $\ldots$
agent 11 11 0 11 1 11 $\ldots$
agent 22 11 1 1 11 1 $\ldots$
agent 3 $a_{3}$ $a_{3}$ $a_{3}$ $b_{3}$ $b_{3}$ $\ldots$
@@
For the deterministic allocation of indivisible chores for (n1)(n-1) agents with additive binary cost functions and one agent with the additive personalized bi-valued cost function, the Adapted-Chores-Picking Algorithm (Algorithm 13) computes a complete allocation that satisfies EF1 and MMS.
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 13.
EF1.
We show that the computed allocation is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after the round $k$, the allocation $A^{k}$ is EF1. We show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is also EF1.
@@
  • If $c_{n}(e_{k+1})=a_{n}$, then let us consider two subcases. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\geq|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is still EF1. For agent $n$, her bundle is EF1 since she has the priority of picking the item with the cost of $a_{n}$, and the number of items in her bundle is no more than that of items with non-zero costs in agent $i_{min}$’s bundle. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|<|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the additive binary cost valuation and the lowest number of items with non-zero costs.
  • If $v_{n}(e_{k+1})=b_{n}$, the let us consider two subcases. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\leq|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the additive binary cost valuation and the lowest number of items with non-zero costs. For any agent $i\in N\setminus\{i_{min}\}$, her bundle remains untouched and is still EF1. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|>|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is trivially EF1. For agent $n$, her bundle is EF since she has the priority of picking the item with the cost of $a_{n}$, and in this case, the extra item in another agent’s bundle must have the cost of $b_{n}$ for agent $n$.
Combining the above cases, it can be seen that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k+1}$ is still EF1.
MMS.
Fix an arbitrary agent $i\in N$ and an arbitrary round $k$. If her bundle is EF, we are done. If not, then let us consider two cases. When $i\in[n-1]$, she has the additive binary valuation function, and by our algorithm description, the number of items with non-zero costs in her bundle is at most one more than that of the other agent’s bundle. Thus, it is not hard to see that her bundle is MMS. When $i=n$, currently, the number of items in her bundle is one more than that of the other agent’s bundle, and the extra item has the cost of $a_{n}$. After removing this item, the cost of her bundle is no more than that of the other agent’s bundle from her perspective, so we can conclude that her bundle is MMS. ∎
@@
Given any $\alpha,\beta\geq 1$, for the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation satisfying
  • $\alpha$-EF1 and $\beta$-min-USC,
  • or MMS and $\beta$-min-USC,
even for two agents where one agent has the additive binary cost function and the other agent has the additive personalized bi-valued cost function.
@@
Without loss of generality, we assume that $c_{1}\in\{0,1\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+1}{2}=\frac{1}{2\epsilon^{2}}+\frac{1}{2}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm outputs a complete allocation that is $\alpha$-EF1 or MMS in the setting of additive binary and additive personalized bi-valued cost functions, and guarantees $\beta$-min-USC for any given $\alpha,\beta\geq 1$. ∎